Given P(E) = 0.46, P(F) = 0.49, and P(E and F) = 0.15, what is P(E or F)?
Answer by ikleyn(53937)

.
Given P(E) = 0.46, P(F) = 0.49, and P(E and F) = 0.15, what is P(E or F)?
~~~~~~~~~


Use a general probability theory identity / formula


    P(E or F) = P(E) + P(F) - P(E and F) = 0.46 + 0.49 - 0.15 = 0.8.    ANSWER

Solved and explained.